# Arithmetic, Commonsense and Hypothetical Reasoning

### 19 Feb 2024 v1

@@ 1 Introduction
However, previous text-to-SQL datasets Yu et al. (2018); Finegan-Dollak et al. (2018); Yaghmazadeh et al. (2017); Iyer et al. (2017); Zhong et al. (2017); Li and Jagadish (2014); Giordani and Moschitti (2012); Popescu et al. (2003); Tang and Mooney (2000); Dahl et al. (1994), have limitations that prevent them from capturing complex reasoning effectively. For example, Spider Yu et al. (2018) purposely excludes questions that would require external knowledge Pan et al. (2023, 2017a, 2017b), like that from common-sense knowledge graphs or mathematical calculations. This exclusion limits Spider’s ability to properly test how well models can handle real-world scenarios, which often require a deeper level of reasoning capabilities.
@@ 2 Reasoning Types
Arithmetic reasoning pertains to the act of resolving mathematical problems through logical and analytical thought processes, involving datatype values Pan and Horrocks (2003, 2005) and arithmetic operators. According to an analysis of SQL queries from practical applications like the Baidu search engine and customer service and data analysis robots by Wang et al. (2020), mathematical calculations account for a significant portion across SQL applications. However, previous high-quality datasets contain very few questions that involve calculations, and such questions are typically auto-generated with simple grammar. In contrast, all question-SQL pairs included in Archer necessitate arithmetic reasoning and are manually annotated to ensure high quality.
@@
Commonsense reasoning refers to the capacity to make logical deductions based on implicit (and possibly uncertain) commonsense knowledge Romero et al. (2019); Arnaout et al. (2022); He et al. (2023); Wan et al. (2021); Wang et al. (2010, 2014); Stoilos et al. (2006); Pan et al. (2005), including, e.g., a broad understanding of how things function in the world. Commonsense knowledge can be useful for both zero-shot learning Chen et al. (2023a, 2021a, 2021b, b); Geng et al. (2023) and model explanations Guan et al. (2024); Chen et al. (2018). Archer includes questions that necessitate models to comprehend the database, infer missing details, and generate logical inferences to create accurate SQL queries. As illustrated in Figure 1, for the question "Which 4-cylinder car needs the most fuel to drive 300 miles? List how many gallons it needs, and its make and model.", the database does not provide an explicit schema about the fuel used to travel 300 miles for each car. It only provides each car’s fuel consumption in MPG. Solving this question requires commonsense knowledge, specifically the understanding of "Fuel used is calculated by dividing distance driven by fuel consumption" to derive the correct SQL.
@@ 7 Related Work
Wang et al. (2020) proposed DuSQL, a Chinese cross-domain text-to-SQL dataset that includes math-related questions. However, DuSQL’s queries and questions are relatively simple due to automatic generation and grammar restrictions. Dou et al. (2022) extended DuSQL with external knowledge in their KnowSQL dataset. Unfortunately, KnowSQL is not publicly unavailable.
@@
In the solution space, there are both LLM based solutions and solutions based on fine-tuned models. The former solutions, such as DIN-SQL Pourreza and Rafiei (2024), tend to perform better in existing text-to-SQL datasets, while the latter ones, particularly FastRAT Vougiouklis et al. (2023), can offer significant improvements on latency, while keeping decent performance. There can be space combining the above two kinds of solutions for Archer, which is a promising direction for future work.

### 25 Feb 2024 v2

@@ 1 Introduction
However, previous text-to-SQL datasets Yu et al. (2018); Finegan-Dollak et al. (2018); Yaghmazadeh et al. (2017); Iyer et al. (2017); Zhong et al. (2017); Li and Jagadish (2014); Giordani and Moschitti (2012); Popescu et al. (2003); Tang and Mooney (2000); Dahl et al. (1994), have limitations that prevent them from capturing complex reasoning effectively. For example, Spider Yu et al. (2018) purposely excludes questions that would require external knowledge, like common-sense inferences or mathematical calculations. This exclusion limits Spider’s ability to properly test how well models can handle real-world scenarios, which often require a deeper level of reasoning capabilities.
@@ 2 Reasoning Types
Arithmetic reasoning pertains to the act of resolving mathematical problems through logical and analytical thought processes. According to an analysis of SQL queries from practical applications like the Baidu search engine and customer service and data analysis robots by Wang et al. (2020), mathematical calculations account for a significant portion across SQL applications. However, previous high-quality datasets contain very few questions that involve calculations, and such questions are typically auto-generated with simple grammar. In contrast, all question-SQL pairs included in Archer necessitate arithmetic reasoning and are manually annotated to ensure high quality.
@@
Commonsense reasoning refers to the capacity to make logical deductions based on implicit knowledge and a broad understanding of how things function in the world. Archer includes questions that necessitate models to comprehend the database, infer missing details, and generate logical inferences to create accurate SQL queries. As illustrated in Figure 1, for the question "Which 4-cylinder car needs the most fuel to drive 300 miles? List how many gallons it needs, and its make and model.", the database does not provide an explicit schema about the fuel used to travel 300 miles for each car. It only provides each car’s fuel consumption in MPG. Solving this question requires commonsense knowledge, specifically the understanding of "Fuel used is calculated by dividing distance driven by fuel consumption" to derive the correct SQL.
@@ 7 Related Work
Then, Wang et al. (2020) proposed DuSQL, a Chinese cross-domain text-to-SQL dataset that includes math-related questions. However, DuSQL’s queries and questions are relatively simple due to automatic generation and grammar restrictions. Dou et al. (2022) extended DuSQL with external knowledge in their KnowSQL dataset. Unfortunately, KnowSQL is not publicly unavailable.